# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

### v2

@@ 4.2 Performance
In this study, we employ our algorithm on the DPM++ 2m solver and UniPC solver. We employ $p_{1}=7$, $p_{2}=1.2$, $stop=8$, $M=12$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images via DPM++ 2m. Additionally, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ are applied on the custom UniPC solver in the eight-step inference when generating 1024 x 1024 images. And for the six-step inference for generating 1024 x 1024-resolution images, the hyperparameters of the UniPC solver are $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$. Meanwhile, for 512 x 512 images, the hyperparameters of the 8-step UniPC solver are $p_{1}=7$, $p_{2}=1.2$, $stop=9$, $M=12$, and $N=8$. And for the customized 5-step UniPC solver for 512 x 512 images, the hyperparameters are $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=5$. Then, the 6-step UniPC solver of the 512 x 512-resolution images has hyperparameters $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=6$.
For the eight-step inference that uses DPM++ 2m to generate 512x512-resolution images, we avoid using free-U. And for the customized DPM++ 2m solver that generates 1024x1024-resolution images, the $t_{aug}$ apply free-U is 8. For all UniPC solver customized to generate 512x512 resolution images, $t_{aug}=4$. And $t_{aug}=6$, $t_{aug}=4$ for 6-step UniPC solver and 8-step UniPC solver customeized for 1024 x 1024-resolution images, respectively. And we only use AFS on the UniPC solver customized for the 1024 x 1024-resolution image. We don’t guarantee we provide the best hyperparameters; we just ensure we provide hyperparameters that can sample images with FID performance comparable to that of the inference result from the 20-step DPM++ solver and the SOTA distillation algorithm.
@@
We report the results of the inference that generate 512 x 512-resolution images in the COCO 2014, COCO 2017, and LAION dataset in table˜2, table˜3, and table˜4. Moreover, the result of the 1024 x 1024 images is given in table˜5, table˜6, and table˜7. Explicitly, the UniPC solver is a 2nd-order multistep solver for eight-step inference and a 1st-order multistep solver for five and six-step inference. The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively. Furthermore, all the synthesized images in table˜2 and table˜5 employs the random 10k COCO caption as a guidance prompt. In table˜3 and table˜6, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance. Furthermore, we use 10k captions in LAION for table˜4 and table˜7. We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2 (Yin et al. (2024a); Chadebec et al. (2025); Lin et al. (2024); Sauer et al. (2024)). Our approach outperforms state-of-the-art models for most of the datasets. However, some of the models (Yin et al. (2024a); Chadebec et al. (2025); Lin et al. (2024)) are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison. The $w$ in the table˜2 and table˜5 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
Solver FID CLIP-Score
(Our) custom DPM++1s(5) $w=7.5$ 19.18 30.92
(Our) custom DPM++1s(6) $w=7.5$ 18.54 31.16
(Our) custom DPM++2m(8) $w=7.5$ 15.92 31.32
(Our) custom DPM++1s(5) $w=5.5$ 17.54 31.0
(Our) custom DPM++1s(6) $w=5.5$ 17.04 31.16
(Our) custom DPM++2m(8) $w=5.5$ 15.7 31.07
(Our) custom UniPC(5) $w=7.5$ 19.35 30.63
(Our) custom UniPC(6) $w=7.5$ 16.41 30.97
(Our) custom UniPC(8) $w=7.5$ 16.00 31.15
(Our) custom UniPC(5) $w=5.5$ 15.19 30.93
(Our) custom UniPC(6) $w=5.5$ 14.72 31.08
(Our) custom UniPC(8) $w=5.5$ 15.26 31.12
Original DPM++2m(20) $w=7.5$ 17.3 30.97
Original DPM++2m(20) $w=5.5$ 18.68 30.97
Flash Diffusion(5) 17.34 30.45
Flash Diffusion(6) 17.94 30.59
Flash Diffusion(8) 19.03 30.37
AMED Pulgin(8) 19.07 31.15
SDXL turbo(4) 23.24 31.66
@@
Solver FID CLIP-Score
(Our) custom DPM++1s(5) $w=5.5$ 24.22 31.06
(Our) custom DPM++1s(6) $w=5.5$ 23.24 31.12
(Our) custom DPM++2m(8) $w=5.5$ 22.40 31.23
(Our) custom DPM++1s(5) $w=7.5$ 25.58 30.93
(Our) custom DPM++1s(6) $w=7.5$ 24.71 31.13
(Our) custom DPM++2m(8) $w=7.5$ 22.35 31.35
(Our) custom UniPC(5) $w=7.5$ 22.70 30.99
(Our) custom UniPC(6) $w=7.5$ 24.02 31.04
(Our) custom UniPC(8) $w=7.5$ 24.55 30.99
(Our) custom UniPC(5) $w=5.5$ 21.31 30.92
(Our) custom UniPC(6) $w=5.5$ 20.82 31.04
(Our) custom UniPC(8) $w=5.5$ 21.71 31.07
Original DPM++2m(20) $w=7.5$ 25.97 31.21
Original DPM++2m(20) $w=5.5$ 23.75 31.10
Flash Diffusion(5) 23.56 30.64
Flash Diffusion(6) 24.16 30.60
Flash Diffusion(8) 25.35 30.55
AMED Pulgin(8) $w=7.5$ 25.50 31.17
SDXL turbo(4) 30.92 31.66
@@
Solver FID CLIP Score
(Our) custom DPM++1s(5) $w=5.5$ 19.61 31.81
(Our) custom DPM++1s(6) $w=5.5$ 18.45 32.10
(Our) custom DPM++2m(8) $w=5.5$ 17.74 32.42
(Our) custom DPM++1s(5) $w=7.5$ 19.88 31.45
(Our) custom DPM++1s(6) $w=7.5$ 18.60 31.97
(Our) custom DPM++2m(8) $w=7.5$ 17.52 32.41
(Our) custom UniPC(5) $w=7.5$ 20.43 30.87
(Our) custom UniPC(6) $w=7.5$ 17.56 31.63
(Our) custom UniPC(8) $w=7.5$ 17.38 32.08
(Our) custom UniPC(5) $w=5.5$ 17.44 31.73
(Our) custom UniPC(6) $w=5.5$ 16.84 32.08
(Our) custom UniPC(8) $w=5.5$ 17.32 32.26
Original DPM++2m(20) $w=7.5$ 18.30 32.67
Original DPM++2m(20) $w=5.5$ 17.33 32.47
Flash Diffusion(5) 15.40 31.60
Flash Diffusion(6) 15.80 31.63
Flash Diffusion(8) 16.87 31.57
AMED Pulgin(8) $w=7.5$ 18.06 32.44
SDXL turbo(4) 22.25 33.02
@@
Solver FID CLIP Score
custom DPM++2m (8) $w=7.5$ 17.84 31.62
custom DPM++2m (8) $w=5.5$ 19.62 31.58
custom UniPC (8) $w=7.5$ 18.18 31.23
custom UniPC (8) $w=5.5$ 17.87 31.40
custom UniPC (6) $w=7.5$ 18.80 31.42
custom UniPC (6) $w=5.5$ 19.07 31.45
Flash DiffusionXL (6) 22.23 31.00
Flash DiffusionXL (8) 23.25 30.89
SDXL-Lightning (8) 21.07 31.11
DMD2 (4) 19.64 31.64
@@
Solver FID CLIP Score
custom DPM++2m (8) $w=7.5$ 24.42 31.71
custom DPM++2m (8) $w=5.5$ 26.04 31.56
custom UniPC (8) $w=7.5$ 24.58 31.21
custom UniPC (8) $w=5.5$ 24.38 31.38
custom UniPC (6) $w=7.5$ 25.29 31.44
custom UniPC (6) $w=5.5$ 25.20 31.42
Flash DiffusionXL (6) 28.74 30.96
Flash DiffusionXL (8) 29.63 30.88
SDXL-Lightning (8) 27.89 31.08
DMD2 (4) 26.19 31.65
@@
Solver FID CLIP Score
custom DPM++2m (8) $w=7.5$ 19.25 32.22
custom DPM++2m (8) $w=5.5$ 20.82 32.07
custom UniPC (8) $w=7.5$ 21.59 31.73
custom UniPC (8) $w=5.5$ 20.41 32.06
custom UniPC (6) $w=7.5$ 22.51 31.82
custom UniPC (6) $w=5.5$ 22.33 31.87
Flash DiffusionXL (6) 20.88 31.64
Flash DiffusionXL (8) 21.30 31.61
SDXL-Lightning (8) 20.45 32.50
DMD2 (4) 16.51 33.28
@@
We also calculate the PRD in a 6-step inference by sampling 10k captions from COCO 2014 and getting the corresponding image in the COCO dataset as a reference. PRD was first used to evaluate GAN. Considering two distributions $Q$ and $P$, the first one is the distribution of the generative images, and the second one is the distribution of the original images, the precision shows how much of $Q$ can be generated by $P$, and the recall exhibits how much $P$ can be generated by $Q$. For example, if $P$ is a subset of MNIST that only contains numbers ranging from 1 to 4, adding 5 and 6 to the generative distribution reduces the precision, while lacking modes like 3 and 4 will reduce the recall rate. Our experiment shows that the recall rate of the flash diffusion in the 6-step is less than that of us. The guidance scale of our sampler is 5.
@@ 5 Conclusion
In this study, we propose a plug-in method to enhance the sampling algorithm. By fully harnessing the potential capacity of the latent diffusion model—such as modifying the skip-connection at the right time, using the property of the decoder, which is a $\beta$-VAE trained with small KL-constraints, ignoring a small amount of noise in the final feature to reschedule the discretizing method, and choosing the proper ODE solver via truncation error analysis—we can sample a high-quality image with limited inference budget and avoid the need for additional training.
@@ 5.1 Limitation and the future work
Firstly, a diffusion model is a robust generative model (Chen et al. (2024)) that can prevent adversarial attacks. Most attacks that hack the latent diffusion model attack its VAE decoder. Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that largely depends on the property of the VAE. Moreover, in future research, we will explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models. Last, the hyperparameters of the ACGN model need further exploration.

### v3

@@ 4.2 Performance
In this study, we employ our algorithm on the DPM++ 2m solver and UniPC solver. We employ $p_{1}=7$, $p_{2}=1.2$, $stop=8$, $M=12$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images via DPM++ 2m. And its $t_{aug}$ is -1 and 8 for 512x512 and 1024x1024 resolutions, respectively. Additionally, $t_{aug}=6$, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ are applied on the custom UniPC solver in the eight-step inference when generating 1024 x 1024 images. And for the six-step inference for generating 1024 x 1024-resolution images, the hyperparameters of the UniPC solver are $t_{aug}=4$, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$. Meanwhile, for 512 x 512 images, the hyperparameters of the 8-step UniPC solver are $t_{aug}=4$, $p_{1}=7$, $p_{2}=1.2$, $stop=9$, $M=12$, and $N=8$. And for the customized 5-step UniPC solver for 512 x 512 images, the hyperparameters are $t_{aug}=4$, $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=5$. Then, the 6-step UniPC solver of the 512 x 512-resolution images has hyperparameters $t_{aug}=4$, $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=6$.
@@
We report the results of the inference that generate images via 10000 captions in the COCO 2014, 5000 captions in COCO 2017, and 10000 captions in the LAION dataset in Table 2, Table 5, Table 3, Table 6, Table 4, and Table 7. Explicitly, the UniPC solver is a 2nd-order multistep solver for eight-step inference and a 1st-order multistep solver for five and six-step inference. The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively. We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2 (Yin et al. (2024a); Chadebec et al. (2025); Lin et al. (2024); Sauer et al. (2024)). Our approach outperforms state-of-the-art models for most of the datasets. However, some of the models (Yin et al. (2024a); Chadebec et al. (2025); Lin et al. (2024)) are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison. The $w$ in the Table 2 and Table 5 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
@@
Solver steps scale use Free-U apply stage FID CLIP-Score
DPM++ 2m 8 7.5 No Never 15.92 31.32
DPM++ 2m 8 5.5 No Never 15.7 31.07
UniPC 5 7.5 Yes 4-th step 19.35 30.63
UniPC 6 7.5 Yes 4-th step 16.41 30.97
UniPC 8 7.5 Yes 4-th step 16.00 31.15
UniPC 5 5.5 Yes 4-th step 15.19 30.93
UniPC 6 5.5 Yes 4-th step 14.72 31.08
UniPC 8 5.5 Yes 4-th step 15.26 31.12
DPM++ 2m 20 7.5 No Never 17.3 30.97
DPM++ 2m 20 5.5 No Never 18.68 30.97
Flash Diffusion 5 None No Never 17.34 30.45
Flash Diffusion 6 None No Never 17.94 30.59
Flash Diffusion 8 None No Never 19.03 30.37
AMED Pulgin 8 7.5 No Never 19.07 31.15
SDXL turbo 4 None No Never 23.24 31.66
@@
Solver steps scale use Free-U apply stage FID CLIP-Score
DPM++ 2m 8 7.5 No Never 22.40 31.23
DPM++ 2m 8 5.5 No Never 22.35 31.35
UniPC 5 7.5 Yes 4-th step 22.70 30.99
UniPC 6 7.5 Yes 4-th step 24.02 31.04
UniPC 8 7.5 Yes 4-th step 24.55 30.99
UniPC 5 5.5 Yes 4-th step 21.31 30.92
UniPC 6 5.5 Yes 4-th step 20.82 31.04
UniPC 8 5.5 Yes 4-th step 21.71 31.07
DPM++ 2m 20 7.5 No Never 25.97 31.21
DPM++ 2m 20 5.5 No Never 23.75 31.10
Flash Diffusion 5 None No Never 23.56 30.64
Flash Diffusion 6 None No Never 24.16 30.60
Flash Diffusion 8 None No Never 25.35 30.55
AMED Pulgin 8 7.5 No Never 25.50 31.17
SDXL turbo 4 None No Never 30.92 31.66
@@
Solver steps scale use Free-U apply stage FID CLIP-Score
DPM++ 2m 8 7.5 No Never 17.74 32.42
DPM++ 2m 8 5.5 No Never 17.52 32.41
UniPC 5 7.5 Yes 4-th step 20.43 30.87
UniPC 6 7.5 Yes 4-th step 17.56 31.63
UniPC 8 7.5 Yes 4-th step 17.38 32.08
UniPC 5 5.5 Yes 4-th step 17.44 31.73
UniPC 6 5.5 Yes 4-th step 16.84 32.08
UniPC 8 5.5 Yes 4-th step 17.32 32.26
DPM++ 2m 20 7.5 No Never 18.30 32.67
DPM++ 2m 20 5.5 No Never 17.33 32.47
Flash Diffusion 5 None No Never 15.40 31.60
Flash Diffusion 6 None No Never 15.80 31.63
Flash Diffusion 8 None No Never 16.87 31.57
AMED Pulgin 8 7.5 No Never 18.06 32.44
SDXL turbo 4 None No Never 22.25 33.02
@@
Solver steps scale use Free-U apply stage FID CLIP-Score
DPM++ 2m 8 7.5 No Never 17.84 31.62
DPM++ 2m 8 5.5 No Never 19.62 31.58
UniPC 8 7.5 Yes 4-th step 18.18 31.23
UniPC 8 5.5 Yes 4-th step 17.87 31.40
UniPC 6 7.5 Yes 4-th step 18.80 31.42
UniPC 6 5.5 Yes 4-th step 19.07 31.45
Flash Diffusion 6 None No Never 22.23 31.00
Flash Diffusion 8 None No Never 23.25 30.89
SDXL-Lightning 8 None No Never 21.07 31.11
DMD2 4 None No Never 19.64 31.64
@@
Solver steps scale use Free-U apply stage FID CLIP-Score
DPM++ 2m 8 7.5 No Never 24.42 31.71
DPM++ 2m 8 5.5 No Never 26.04 31.56
UniPC 8 7.5 Yes 4-th step 24.58 31.21
UniPC 8 5.5 Yes 4-th step 24.38 31.38
UniPC 6 7.5 Yes 4-th step 25.29 31.44
UniPC 6 5.5 Yes 4-th step 25.20 31.42
Flash Diffusion 6 None No Never 28.74 30.96
Flash Diffusion 8 None No Never 29.63 30.88
SDXL-Lightning 8 None No Never 27.89 31.08
DMD2 4 None No Never 26.19 31.65
@@
Solver steps scale use Free-U apply stage FID CLIP-Score
DPM++ 2m 8 7.5 No Never 19.25 32.22
DPM++ 2m 8 5.5 No Never 20.82 32.07
UniPC 8 7.5 Yes 4-th step 21.59 31.73
UniPC 8 5.5 Yes 4-th step 20.41 32.06
UniPC 6 7.5 Yes 4-th step 22.51 31.82
UniPC 6 5.5 Yes 4-th step 22.33 31.87
Flash Diffusion 6 None No Never 20.88 31.64
Flash Diffusion 8 None No Never 21.30 31.61
SDXL-Lightning 8 None No Never 20.45 32.50
DMD2 4 None No Never 16.51 33.28
We also calculate the PRD in a 6-step inference by sampling 10k captions from COCO 2014 and getting the corresponding image in the COCO dataset as a reference. Our experiment shows that the generative diversity of the flash diffusion in the 6-step is less than that of us. The guidance scale of our sampler is 5.
@@ 5 Conclusion
In this study, we propose a plug-in method to enhance the sampling algorithm. By fully harnessing the potential capacity of the latent diffusion model—such as modifying the skip-connection at the right time, using the property of the decoder, ignoring a small amount of noise in the final feature to reschedule the discretizing method, and choosing the proper ODE solver via truncation error analysis—we can sample a high-quality image with limited inference budget and avoid the need for additional training.
@@ 5.1 Limitation and the future work
In future research, we will explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
Appendix A Proof of Theorem 3.1
Firstly, consider the well-known probably approximately correct (PAC)(Hafez-Kolahi et al., 2020) deduction from Hoeffding’s inequality, we have:
$\Delta(s)=E_{X,Y}[\mathcal{L}(f^{s}(X),Y)]-\frac{1}{m}\Sigma_{i=1}^{m}\mathcal{L}(f^{s}(x_{i}),y_{i})\leq\sqrt{\frac{\log{|\mathcal{H}|+\log{\frac{2}{\sigma}}}}{2o}},$ (13)
with probability approaching 1, where $o$ is a given constant for each hypothesis space
And input $X^{(m)}=\{x_{1},x_{2},...,x_{m}\}$, in which sample $x_{i}$ samples i.i.d from $\mathcal{X}$, is a sample from the $\mathcal{X}^{(m)}$. The $|\mathcal{H}|\leq 2^{|\mathcal{X}^{(m)}|}$. Then, for the typical set $\mathcal{A}_{\epsilon}^{(m)}$ of the $\mathcal{X}^{(m)}$, that the probability of $X^{(m)}$ in $\mathcal{A}_{\epsilon}^{(m)}$ is bigger than $1-\epsilon$, and its cardinality $|\mathcal{A}_{\epsilon}^{(m)}|\leq 2^{m(H(x)+\epsilon)}$, then, when the batch size itself is large and the $\epsilon$ is small enough, we have:
$\Delta(s)\leq\sqrt{\frac{2^{mH(X)}+\log{\frac{2}{\sigma}}}{2o}}$ (14)
Where $H(X)$ is the entropy of the $X$. Then, seperate the neural network into two parts: decoder $g^{s}$ and encoder $\phi^{s}$, and $f^{s}=g^{s}\circ\phi^{s}$, where we use $\circ$ to denote the composition of functions. And we use $h^{s}_{i}$ to denote the $i$-th layer of the $n$-layer neural network. Explicilty, $\phi^{s}_{l}=h^{s}_{l}\circ h^{s}_{l-1}\circ\dots\circ h^{s}_{1}$ and $g^{s}_{l}=h^{s}_{n}\circ h^{s}_{n-1}\circ\dots\circ h^{s}_{l+1}$. Then, the cardinality of the hypothesis space of the decoder $g^{s}_{l}$ is constrained by the number of input features $z_{l}^{s}$. And for the hypothesis space $\mathcal{H}_{g}$ with input $Z_{l}^{s}$:
$\Delta(s)\leq\sqrt{\frac{2^{mH(Z_{l}^{s})}+\log{\frac{2}{\sigma}}}{2o}}$ (15)
And then we consider the conditional typical set given by the $p(Z_{l}^{s}|Y)$. When the $Y$ is given, the cardinality of the typical set of $\{{Z_{l}^{s}}^{(1)},{Z_{l}^{s}}^{(1)},\dots,{Z_{l}^{s}}^{(m)}\}$ is $2^{mH(Z_{l}^{s}|Y)}$. Hence for each decoder $g^{s}_{l}$ with input $Z_{l}^{s}$ that having a mutual information constrain $I(Z_{l}^{s};Y)$, the total possible middle feature set is equal to:
$|\mathcal{H}_{g}|=\frac{2^{H(Z_{l}^{s})}}{2^{H(Z_{l}^{s}|Y)}}=2^{I(Z_{l}^{s};Y)}$ (16)
Then, for each decoder $g^{s}_{l}$, we have
$\Delta(s_{g})\leq\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+log\frac{2}{\sigma}}{2o}}$ (17)
Meanwhile, for each encoder $\phi_{l}^{s}$, we use $I(\phi_{l}^{s};S)$ to constrain the overfitting. Explicitly, the cardinality of the hypothesis space $\mathcal{H_{\phi}}$ of the encoder is $2^{I(\phi_{l}^{s};S)}$ (Kawaguchi et al., 2023). And then, we mulitiply the cardinality of two hypothesis spaces $\mathcal{H}_{g}$ and $\mathcal{H}_{\phi}$, to get the hypothesis space of $f_{l}^{s}$ whose the decoder is constrained with mutual information, and the encoder is in the $\mathcal{H_{\phi}}$, we have:
$\Delta(s_{l})\leq\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};S)}+log\frac{2}{\sigma}}{2o}}$ (18)
And the $\Delta(s)$, when each encoder is constrained by a mutual information $I(\phi_{l}^{s};S)$ to prevent overfiting, is:
$\Delta(s)\leq\underset{l}{\arg\min}\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};S)}+log\frac{2}{\sigma}}{2o}}$ (19)
Then, if we ignore the order of output in a batch of validation data, we have
$\Delta(s)\leq\underset{l}{\arg\min}\sqrt{\frac{2^{mI(Y;Z_{l}^{s})}+{I(\phi_{l}^{s};S)}+log\frac{2}{\sigma}}{2o*A_{m}^{m}}}$ (20)
and A is the arrangement number, which means the mutual information gives a bound on the generalization error. $\Box$
Appendix B Proof of Proposition 3.3
For the generalization error $E_{X,Y}[{\mathcal{L}(f^{s}(X),Y)}]$, we have
$\mathcal{L}(f^{s}(x),y)$ $=(f^{s}(x)-y)^{2}$ (21)
$=(f^{s}(x)-E[y|x]+E[y|x]-y)^{2}$ (22)
$=(f^{s}(x)-E[y|x])^{2}+(E[y|x]-y)^{2}+2(f^{s}(x)-E[y|x])(E[y|x]-y).$ (23)
Recall that
$E_{X,Y}[\mathcal{L}]$ $=\int_{x}\int_{y}\mathcal{L}(f^{s}(x),y)p(x,y)dxdy.$ (24)
By (24), we compute first and third term on the right hand side of (23):
$\int_{x}\int_{y}(f^{s}(x)-E[y|x])^{2}p(x,y)dxdy$ $=\int_{x}(f^{s}(x)-E[y|x])^{2}p(x)dx,$ (25)
$\int_{x}\int_{y}2(f^{s}(x)-E[y|x])(E[y|x]-y)dxdy$ $=0.$ (26)
Therefore, we have
$E_{X,Y}[\mathcal{L}]$ $=\int_{x}(f^{s}(x)-E[y|x])^{2}p(x)dx+\int_{x}\int_{y}(E[y|x]-y)^{2}p(x,y)dxdy,$ (28)
where
$\int_{x}(f^{s}(x)-E[y|x])^{2}p(x)dx$ $=E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]+E_{X,Y}[(\hat{f}(X)-E[y|x])^{2}]$ (29)
where $E_{X}[(f^{s}(X)-\hat{f}(X))^{2}]$ denote the distance of the output from a particular training neural network $f^{s}$ and the average output $\hat{f}$. Note that
$\int_{x}\int_{y}(E[y|x]-y)^{2}p(x,y)dxdy+E_{X,Y}[(\hat{f}(X)-E[y|x])^{2}]=E_{X,Y}[(\hat{f}(X)-y)^{2}].$ (30)
Combining (28), (29) and (30), we derive the final result. $\Box$